# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

### v3

@@ 6 Conclusions
We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property. While this paper focused on a rigorous exposition of the AEP property, it opens up many questions for further study. An important open problem is to design practical algorithms for AI text detection and dataset inference using the typical set concept. Tighter non-asymptotic bounds on deviations that improve on (18) will significantly benefit practical applications by providing better performance guarantees. More fundamentally, Proposition 3.3 leads naturally to the question of how much do the typical sets of different models overlap. The theory presented in this paper is neutral on this question which must be answered empirically and both possible answers have interesting implications. No overlap between typical sets means that each model can be associated with a unique statistical signature. On the other hand, independently trained models converging on a common set of typical strings, would be a notable linguistic phenomenon, though it is hard to speculate on what this might mean. The AEP is also just one example of the broader observation that generative AI models are stochastic machines subject to statistical laws and the discovery of other statistical regularities in the outputs of generative models is an important topic for future work.
Appendix A A Simple Markov Language Model
We now illustrate the ideas from Section 3 for a simple Markov language model, $M_{1}$ with a binary token space consisting of only two tokens $\mathcal{T}=\{t_{1},t_{2}\}$ and a parameter $\rho\in(0,1)$. The user prompt ${\mathbf{X}}$ consists of a single token which we can set to be $t_{1}$ without loss of generality: ${\mathbf{X}}=[t_{1}]$. The model $M_{1}$ generates a text string of length $N$ following a simple probabilistic algorithm: the $n$’th token $Y_{n}$ is set equal to the $n-1$’th token with probability $\rho$. The prompt serves as the initial token i.e. $Y_{0}\doteq t_{1}$. Let $h_{b}(\rho)\doteq-\rho\log_{2}\rho-(1-\rho)\log_{2}(1-\rho)$ denote the binary entropy function with argument $\rho$.
It is easy to see that $H(p_{n})\equiv h_{b}(\rho),~{}\forall n$ and $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(\rho)$ for any text string ${\mathbf{Y}}$. For $\rho=0.9$, $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(0.9)\approx 0.47$
Example A.1.
Consider the string ${\mathbf{Y}}_{1}$ with all tokens $Y_{n}=t_{1},~{}\forall n=1\dots N$. The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(0.9)\approx 0.15$.
Example A.2.
Consider the string ${\mathbf{Y}}_{2}$ with alternating token values: ${\mathbf{Y}}_{2}\equiv[t_{2},t_{1},t_{2},t_{1},t_{2},t_{1}\dots]$. The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(1-\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(0.1)\approx 3.3$.
Note that ${\mathbf{Y}}_{1}$ and ${\mathbf{Y}}_{2}$ are the most likely and least likely string of length $N$ respectively for $\rho=0.9$ (or indeed for any $\rho>0.5$). However, it is clear that neither string is typical for any reasonable $\epsilon$ with the perplexity of ${\mathbf{Y}}_{1}$ being abnormally low and that of ${\mathbf{Y}}_{2}$ abnormally high respectively.
Example A.3.
Consider the string ${\mathbf{Y}}_{3}$ created by flipping a fair coin repeatedly to generate a vector whose elements are iid randomly selected with $\Pr(t_{1})=\Pr(t_{2})=0.5$. Since $Y_{3}$ is a random string, its perplexity $l_{M_{1}}({\mathbf{Y}}_{3})$ is also a random variable. However, for long strings i.e. $N\gg 1$, with high probability, $l_{M_{1}}({\mathbf{Y}}_{3})\approx-0.5\log_{2}\rho-0.5\log_{2}(1-\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{3})\approx 1.74$ which is incidentally the arithmetic mean of $l_{M_{1}}({\mathbf{Y}}_{1})$ and $l_{M_{1}}({\mathbf{Y}}_{2})$.
Clearly, with high probability, the string ${\mathbf{Y}}_{3}$ is not typical. Indeed, from these examples, we can see that typical strings for the model $M_{1}$ are precisely those that contain close to $(1-\rho)N$ token ’flips’ i.e. instances when $Y_{n+1}\neq Y_{n}$.
Appendix B A Generalized Law of Large Numbers for Random Variables with Arbitrary Dependence
@@
B.1 Conditional Means and Sample Means for Sequences of Dependent Random Variables
@@
$\mu_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)g_{n}({\mathbf{Y}}_{n})\equiv\operatorname*{{E}}\left[Z_{n}|{\mathbf{Y}}_{n-1}\right]$ (26)
$V_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)\left(g_{n}({\mathbf{Y}}_{n})-\mu_{n}^{({\mathbf{Y}})}\right)^{2}\equiv\operatorname*{{E}}\left[Z_{n}^{2}|{\mathbf{Y}}_{n-1}\right]-\left(\mu_{n}^{({\mathbf{Y}})}\right)^{2}$ (27)
@@
Clearly, $\operatorname*{{E}}\left[Z_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{n}^{({\mathbf{Y}})}\right]$ and $\operatorname*{{E}}\left[S_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{S_{n}}^{({\mathbf{Y}})}\right]$.
Lemma B.1.
@@
$=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]-2\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]$ (33)
@@
Remark. An alternative way to arrive at (35) is to treat $\mu_{l+1}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{l+1}\big{|}{\mathbf{Y}}_{l}\right]$ as an estimate of $Z_{l+1}$ given ${\mathbf{Y}}_{l}$, and then use the famous Orthogonality Principle to argue that the estimation error $\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)$ must be uncorrelated with any function of ${\mathbf{Y}}_{l}$. More formally, $\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)$ is $\mathcal{F}_{l}$-measurable and is therefore orthogonal to $\left(Z_{l+1}-\operatorname*{{E}}\left[Z_{l+1}|\mathcal{F}_{l}\right]\right)$.
@@
Using Lemma B.1, we can write:
@@
Lemma B.2 states that the sample means of the two random sequences $\{Z_{n}\}$ and $\{\mu^{({\mathbf{Y}})}_{n}\}$ must be close to each other asymptotically. Since we have made very few assumptions about the random variables involved, we cannot assume that either of the sequences $\{Z_{n}\}$ or $\{\mu^{({\mathbf{Y}})}_{n}\}$ themselves converge in any sense to an asymptotic limit. Neither can we assume that their sample means separately converge to a limit.
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{m=1}^{n}Z_{n}$ $\to_{p}\mu^{({\mathbf{Y}})}~{}~{}\mathrm{or}~{}\lim_{n\to\infty}\Pr\left(\left|\frac{1}{n}S_{n}-\mu^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~{}\forall\epsilon>0$ (39)
@@
Combining Lemma B.2 with the observation $\frac{1}{n}\mu_{S_{n}}^{({\mathbf{Y}})}\rightarrow\mu^{({\mathbf{Y}})}$ completes the proof. ∎
B.2 Discussion
Proposition B.1 is remarkable because unlike traditional versions of the LLN, it makes no assumptions whatsoever about the dependence or correlation between the symbols $Y_{n}$. It does not even require for instance that the correlation between $Y_{1}$ and $Y_{n}$ should decrease as $n$ becomes large.
The two extreme cases of strong dependence and complete independence for the symbols $Y_{n}$ turn out to be two simple special cases for our proposed LLN:
  • If the symbols $Y_{n}$ are independent, Proposition B.1 simply reduces to the classical weak LLN.
  • If the symbols are perfectly correlated, the $Z_{m}$ become deterministic functions of the first token, and the LLN then becomes a trivial observation.
Appendix C Supplemental Material
C.1 Additional Experimental Context and Results
In this work, all experiments were ran on a high-performance computing server running Linux (Ubuntu). AMD EPYC 7413 24-Core Processor, 48 logical cores, 96 GB RAM, four NVIDIA A100-SXM4-80GB GPUs. On this setup, each run of a text generation (i.e., running GPT-2 to generate Nmax = 512 tokens, Llama-8B to generate Nmax = 1024 tokens and Nmax = 10240 tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds to minutes, with each experiment taking about 90 minutes to run in the case of 10,240 steps. We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.3.
Figure 5 shows additional results similar to Fig. 2 where we naively varied the seed value of the random number generator from $209-214$ for Llama-8B. We present these to show that Fig. 2’s results are representative of the general case and were not cherry-picked. In all cases, $K=100$.
(a) Seed=209.
(b) Seed=210.
(c) Seed=211.
(d) Seed=212.
(e) Seed=213.
(f) Seed=214.
Figure 5: Log-perplexity and empirical entropy for Llama 3.1 of strings generated by the same model with top-100 sampling.

### v4

@@ 7 Conclusions
We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property. We used this property to extend the information theoretic concept of typical set to language models. The main take-aways are as follows:
  1. 1.
    The classical AEP and the concept of typical sets can be generalized to apply to the outputs of LLMs.
  2. 2.
    These typical sets can be shown to be a vanishingly small subset of possible ”good” outputs under very general conditions and this means that LLMs are strongly constrained in what they can generate.
  3. 3.
    Experimentally we see that the asymptotic behavior in the AEP can be observed for LLMs for moderate length texts (a few hundred tokens), and thus may be practically relevant to detecting texts of abnormally high or low perplexity.
  4. 4.
    Synthetic texts from LLMs are random processes with known statistics (unlike natural text), and discovering other laws (e.g. Central Limit Theorems) that apply to LLM outputs is an open research question.
Appendix A A Generalized Law of Large Numbers for Random Variables with Arbitrary Dependence
@@
A.1 Conditional Means and Sample Means for Sequences of Dependent Random Variables
@@
$\mu_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)g_{n}\left([{\mathbf{Y}}_{n-1}~y]\right)\equiv\operatorname*{{E}}\left[Z_{n}|{\mathbf{Y}}_{n-1}\right]$ (28)
$V_{n}^{({\mathbf{Y}})}$ $\doteq\sum_{y\in\mathcal{Y}}p^{({\mathbf{Y}})}_{n}(y)\left(g_{n}\left([{\mathbf{Y}}_{n-1}~y]\right)-\mu_{n}^{({\mathbf{Y}})}\right)^{2}\equiv\operatorname*{{E}}\left[Z_{n}^{2}|{\mathbf{Y}}_{n-1}\right]-\left(\mu_{n}^{({\mathbf{Y}})}\right)^{2}$ (29)
@@
Clearly, $\operatorname*{{E}}\left[Z_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{n}^{({\mathbf{Y}})}\right]$ and $\operatorname*{{E}}\left[S_{n}\right]\equiv\operatorname*{{E}}\left[\mu_{S_{n}}^{({\mathbf{Y}})}\right]$. The key idea is that the sequence $Z_{m}-\mu_{m}^{({\mathbf{Y}})}$ is a martingale difference sequence, and $S_{n}-\mu_{S_{n}}^{({\mathbf{Y}})}$ is a martingale for any arbitrary distribution of the $Z_{m}$’s.
Lemma A.1.
@@
$=V_{S_{l}}+\operatorname*{{E}}\left[V_{l+1}^{({\mathbf{Y}})}\right]+2\operatorname*{{E}}\left[\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)\right]$ (35)
@@
Remark. An alternative way to arrive at (37) is to treat $\mu_{l+1}^{({\mathbf{Y}})}\equiv\operatorname*{{E}}\left[Z_{l+1}\big{|}{\mathbf{Y}}_{l}\right]$ as an estimate of $Z_{l+1}$ given ${\mathbf{Y}}_{l}$, and then use the Orthogonality Principle to argue that the estimation error $\left(Z_{l+1}-\mu_{l+1}^{({\mathbf{Y}})}\right)$ must be uncorrelated with any function of ${\mathbf{Y}}_{l}$. More formally, $\left(S_{l}-\mu_{S_{l}}^{({\mathbf{Y}})}\right)$ is $\mathcal{F}_{l}$-measurable and is therefore orthogonal to $\left(Z_{l+1}-\operatorname*{{E}}\left[Z_{l+1}|\mathcal{F}_{l}\right]\right)$.
@@
Using Lemma A.1, we can write:
@@
Lemma A.2 states that the sample means of the two random sequences $\{Z_{n}\}$ and $\{\mu^{({\mathbf{Y}})}_{n}\}$ must be close to each other asymptotically. Since we have made very few assumptions about the random variables involved, we cannot assume that either of the sequences $\{Z_{n}\}$ or $\{\mu^{({\mathbf{Y}})}_{n}\}$ themselves converge in any sense to an asymptotic limit. Neither can we assume that their sample means separately converge to a limit.
@@
$\frac{1}{n}\sum_{m=1}^{n}Z_{n}$ $~\begin{subarray}{c}\rightarrow\\ p\end{subarray}~\mu^{({\mathbf{Y}})}~~\mathrm{or}~\lim_{n\to\infty}\Pr\left(\left|\frac{1}{n}S_{n}-\mu^{({\mathbf{Y}})}\right|>\epsilon\right)=0,~\forall\epsilon>0$ (41)
@@
Combining Lemma A.2 with the observation $\frac{1}{n}\mu_{S_{n}}^{({\mathbf{Y}})}\rightarrow\mu^{({\mathbf{Y}})}$ completes the proof. ∎